OpenStudy (anonymous):

I forgot how to do this type of problem. Input in standard form the equation of the given line.The line with m = -1 /2 and b = 1 I forgot how to do this type of problem. Input in standard form the equation of the given line.The line with m = -1 /2 and b = 1 @Mathematics

OpenStudy (anonymous):

y=mx+b y=-1/2x+1

OpenStudy (anonymous):

y=-1/2x+1 where m is the slope and b is the y-intercept

OpenStudy (shaik0124):

eqn of line in standrd form is y=mx+b,substitute m and b in this u will get the answer